OpenStudy (hapahearthero):

for f(x)=3x^4-12x^3+6 a) find f'(x) b) at x=1, the slope of the graph of f is ____ ? c) at x=1, the equation of the tangent line is y=____? d) the tangent line is horizontal at x=___? someone please explain to me how to solve this.

OpenStudy (sweetburger):

Ok show me what you have done so far.

OpenStudy (hapahearthero):

\[f(x)= 3x ^{4}-12x ^{3}+6 \] is a better view of the question

OpenStudy (hapahearthero):

\[3(x ^{4}-4x ^{3}+2) =\]

OpenStudy (hapahearthero):

I am not really sure where to go from there

OpenStudy (hapahearthero):

for the f'(x) i got 12x^2(x-3)

OpenStudy (sweetburger):

Well here it doesn't necessarily matter if you factor out a 3 or not. I would actually just not factor the 3 out at all. so \[f(x)=3x^4-12x^3+6\] now just differentiate each term of the polynomial. Which is shown \[nx^\left( n-1 \right)\]

OpenStudy (hapahearthero):

so 12x^3-36x^2

OpenStudy (hapahearthero):

how do you figure the slope of the graph at a specific point

OpenStudy (sweetburger):

yes so \[f'(x)=12x^3-36x^2\] or how you factored it \[f'(x)=12x^2(x-3)\]

OpenStudy (sweetburger):

Now what is the slope of the tangent line at x=1?

OpenStudy (hapahearthero):

yes it says at x=1 what is the slope of f

OpenStudy (hapahearthero):

do I graph this to help me?

OpenStudy (sweetburger):

no you solve this analytically

OpenStudy (sweetburger):

find \[f'(1)=12x^3-36x^2\]

OpenStudy (hapahearthero):

-24

OpenStudy (hapahearthero):

then what is the equation of the tangent line at x=1

OpenStudy (hapahearthero):

ah i see where that inputting the one comes in. that makes sense

OpenStudy (sweetburger):

Ok so you have to use point-slope form. You are given the slope of the line at x=1 you are given the point x=1 now all you must find is the corresponding y value by plugging in x=1 into the original function.

OpenStudy (hapahearthero):

ah -3 right

OpenStudy (sweetburger):

So (1,-3) m=-24 into point slope

OpenStudy (hapahearthero):

Oh i thought you mean the original function like the given question... ah

OpenStudy (sweetburger):

Wait did i confuse you on a step?

OpenStudy (hapahearthero):

we would be using y=mx+b correct?

OpenStudy (sweetburger):

Well you could.

OpenStudy (hapahearthero):

so b is 21

OpenStudy (sweetburger):

yep

OpenStudy (hapahearthero):

awesome thank you!!! great help

OpenStudy (sweetburger):

np glad to help :)

OpenStudy (sweetburger):

Also the tangent line is horizontal when the f'(x) = 0 so \[12x^3-36x^2=0\]